INAF queue observing, 2016 November 8 - 9

Observer: F. Cusano, A. Rossi
Telescope Operator: G. Bechetti
Support Astronomer: D. Thompson on call
Instrument: LUCI


At the sunset the sky is clear
LUCI2 still not available
We use LUCI1.

Half Night duration (12\xB0 el twilight to midnight) 18:30-0:00 ( 01:30-7.00 UT) = 5h 30m

Observing time: 4h10m (76%)

Weather time loss: 40m (2%)

technical time loss 40m (2%)

Data Summary

Project_RUN INST Objects Filter ExpTime Notes
Skyflat LUCI   Ks    
Moscadelli J LUCI   Ks, H2 5minKs, 1h25mH2 ONLY LUCI 1
Moscadelli C LUCI   Ks, H2 5minKs, 40m H2 ONLY LUCI 1
dark LUCI darks blind 6ditx25ndit,2.74x18,2.74x10,6x10 ok for Moscadelli
calibration LUCI flats GR200,zJ+HK Ls0.75   ok for Antoniucci
calibration LUCI arcs GR200,zJ+HK Ls0.75   ok for Antoniucci


0:40 taking sky flat with LUCI1 in Ks
  • luci1.20161108.0001-0011 not good
  • luci1.20161108.0012-31 should be ok, counts between 3-11K

1:30 slewing to Piconcelli S program ,starting with the telluric
  • seeing 1.5"
  • preset cancelled at 1:47 problem with the GCS
  • sent preset again at 1:49 tech loss 20min
  • problem with the scripts wrong filter cross match

2:13 slew to Moscadelli program
  • wind near the limits of 20 m/s

2:15 starting with standard field FS149 in H2
  • telescope is not guiding, the script has no guide star luci1.20161108.0040-0044 Ks

2:20 standard FS149 in Ks luci1.20161108.0045-0049 Ks

2:25 slew to science field Moscadelli J, starting with Ks
  • seeing 1" on guider at 2:30 luci1.20161108.0050-0055 Ks

2:36 executing H2 of Moscadelli J luci1.20161108.0056-0083 H2 (would be 80 but we added other 3 exposures later, see below)
  • seeing 0.9" on the guider at 2:40
  • seeing 0.7" on images at 3:02
  • seeing 1.1" on guider at 3:25
  • wind rear 22 m/s, seeing increasing now (3:32) at 1.5"
  • seeing 1.1" on images at 3:36

3:55 we can not point to Iovino due to wind restrictions

04:00 tech problem : 10 loss

4:10 executing again H2 of Moscadelli J: just 3 exposures seeing 0.7" on guider at 4:15

4:20 we go again to the stadard FS151 without guide star, Ks +H2
  • luci1.20161108.0084-0093

4:36 slew to science field Moscadelli C, starting with Ks seeing 1.1" on DIMM
  • luci1.20161108.0094-0099
  • seeing 0.8" on guider at 4:38
  • luci1.20161108.0100-0115 H2 images , start at 4:47
  • seeing 0.8" on guider at 4:50
  • seeing 0.7" on guider at 5:05

5:40 tech loss 10 min GUI problem

5:35 slew again to science field Moscadelli J, starting with H2
  • luci1.20161108.00116-0125
  • seeing 0.9" on guider at 5:55
  • seeing 0.7" on guider at 6:05

6:20 close due to wind tec loss 40m

-- darks for Moscadelli

7:00 INAF night is over

8:27 startin calibration for Antoniucci
  • flat Halo2, configuration 200H+K, filter zJ luci1.20161109.0184-188
  • flat Halo2, configuration 200H+K, filter HK luci1.20161109.0190-194

  • arclamp NEON+ARGON, configuration 200H+K, filter zJ luci1.20161109.0195-205
  • arclamp NEON+ARGON+XENON, configuration 200H+K, filter HK luci1.20161109.0207-228

-- %USERSIG{FeliceCusano - 2016-11-09}%


Topic revision: r7 - 09 Nov 2016, AndreaRossi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback